OpenStudy (anonymous):

At a high school championship game 1200 tickets were sold. Student tickets cost $1.50 each and adult tickets cost $5.00 each. The total revenue collected for the game was $3200. How many student tickets were sold? How many adult tickets were sold? URGENT. I really do not know how to do this and on top of that I have to use all 3 solving methods, elimination, graphing, and substitution. please help 😩

OpenStudy (anonymous):

let adults=x 5x+1.50(1200-x)=3200 solve for x or let adults=x and students=y x+y=1200 5x+1.50 y=3200 find x and y

OpenStudy (anonymous):

for your options use second method

OpenStudy (anonymous):

for elimination 5x+5y=6000 subtract and find y and then find x for substitution y=1200-x 5x+1.50(1200-x)=3200 find x and then y